### v1

# Neural reproducing kernel Banach spaces and representer theorems for deep networks

Abstract.
Studying the function spaces defined by neural networks helps to understand the corresponding learning models and their inductive bias. While in some limits neural networks correspond to function spaces that are reproducing kernel Hilbert spaces, these regimes do not capture the properties of the networks used in practice. In contrast, in this paper we show that deep neural networks define suitable reproducing kernel Banach spaces. These spaces are equipped with norms that enforce a form of sparsity, enabling them to adapt to potential latent structures within the input data and their representations. In particular, leveraging the theory of reproducing kernel Banach spaces, combined with variational results, we derive representer theorems that justify the finite architectures commonly employed in applications. Our study extends analogous results for shallow networks and can be seen as a step towards considering more practically plausible neural architectures.
1. Introduction
Neural networks define functions by composing linear and nonlinear maps in a multi-layer (deep) architecture. While easy to implement, the corresponding models are hard to analyze since they are nonlinearly parameterized. Understanding the properties of the function spaces defined by different neural network architectures can give insights in the corresponding learning models. Further, it can provide indications on the underlying inductive bias, namely what functions can be approximated and learned efficiently by a given class of networks.
In some overparametrized regimes, neural networks can be seen to define Hilbert spaces of functions and in particular reproducing kernel Hilbert spaces (RKHS) [1]. For instance, a classical observation is that shallow networks with infinitely many random units correspond to RKHS, with reproducing kernels depending on the considered nonlinearity [15]. This regime, also known as the Gaussian Process (GP) limit, has connections with models such as random features [19].The limits of more complex, possibly non shallow architectures can also be derived and characterized in terms of RKHS, see e.g. [12]. Another infinite-width limit in which neural networks are described by RKHS is the so-called lazy training regime [9]. In this limit, the network weights evolve little during the optimization and can be well approximated by a linear approximation around a random initialization. In this case as well, the corresponding function spaces are RKHS, and the associated kernel is called neural tangent kernel (NTK) [13]. Again, neural tangent kernels and corresponding RKHS can be derived for a variety of architectures, see e.g. [5].
In fact, neither the above GP/random features limit nor the NTK/lazy training regime seem to capture key aspects of neural networks models [11, 4]. Results for shallow networks suggest that neural networks might favor functions with small norms that are not Hilbertian but rather associated to Banach spaces [2]. In turn, representer theorems associated to such norms allow to characterize finite-width networks commonly used in practice [22, 16, 24, 17, 18]. These observations have sparked considerable interest in understanding the Banach spaces associated to neural networks. Notably, one possibility is to consider extensions of classical splines [28, 27, 26, 25]. Another possibility is to consider reproducing kernel Banach spaces (RKBS) [30, 14], see e.g. [3] and references therein.
In this paper, we develop the latter approach tackling the extension from shallow to deep networks. The study of Banach spaces associated to deep architectures and corresponding representer theorems was started in [18], where deep architectures with ReLU activations and finite rank constraints at each layer are considered. The latter requirement is not natural and is mainly due to technical reasons. Indeed, the finite rank constraint allows for the construction of layers as concatenation of real valued functions studied for shallow networks.
In our study, we propose an approach which allows to consider more general activations and especially to avoid the finite rank constraints. This requires more substantial developments using vector Radon measures to address the challenges posed by potentially infinite-dimensional hidden layers. Our first contribution is to define a reproducing kernel Banach space which describes an infinite-width limit of a deep neural network with an associated norm promoting sparsity. We dub such a space a neural RKBS. Then, we provide a representer theorem for a large class of nonlinearities that shows how optimal neural networks minimizing empirical objectives can be taken to have a finite structure at every layer. This result extends analogous results for shallow networks. It implies that commonly used networks are optimal in the sense that they are solutions of a suitable variational problem.
The rest of the paper is organized as follows. In Section 2 we review the main technical ingredients of our construction, namely reproducing kernel Banach spaces and vector Radon measures. In Section 3 we introduce several notions of RKBS to model functional properties of neural networks, leading to the construction of deep and neural RKBS. In Section 4 we prove our represent theorems on deep and neural RKBS. In Appendix A we collect variational results and extreme point characterizations used to prove our representer theorems.
Table 1. Notation
symbol definition symbol definition
$B(X,Y)$ bounded linear maps $X\to Y$ $\mathcal{M}(\Theta)$ $\mathcal{M}(\Theta,\mathbb{R})$
$X^{\prime}$ continuous dual of $X$ $\|\mu\|_{\operatorname{TV}}$ total variation norm of measure $\mu$
,${}_{X}\langle\cdot,\cdot\rangle_{X^{\prime}}$ pairing on $X,X^{\prime}$ $\delta_{\theta}$ Dirac delta at $\theta$
$\langle\cdot,\cdot\rangle_{X}$ inner product on $X$ $\mathcal{C}_{0}(\Theta,Y)$ continuous functions $\Theta\to Y$ vanishing at $\infty$
$\|\cdot\|_{X}$ norm on $X$ $\mathcal{C}_{0}(\Theta)$ $\mathcal{C}_{0}(\Theta,\mathbb{R})$
$\mathcal{B}(\Theta)$ Borel $\sigma$-algebra on $\Theta$ $B_{X}(r)$ ball on $X$ of radius $r$
$\mathcal{M}(\Theta,Y)$ vector measures on $\Theta$ with values in $Y$ $\operatorname{Ext}(Q)$ extremal points of $Q$
2. Preliminaries
In this section, we review the notion of an architecture and building blocks of deep neural networks. Our perspective is to cast the construction of neural networks in the theory of reproducing kernel Banach spaces, as functions parametrized by vector measures.
Neural networks
We start by setting up some notation, considering to the case of fully connected, feed-forward architectures.
Definition 2.1 (Fully connected feed-forward neural network).
Let $\sigma:\mathbb{R}\to\mathbb{R}$ be a (nonlinear) function, $L\geq 1$ an integer and $d=d_{0},d_{1},\dots,d_{L},d_{L+1}=p\geq 1$ a family of integers. A (fully connected feed-forward) neural network from $\mathbb{R}^{d}$ into $\mathbb{R}^{p}$ with activation function $\sigma$, depth $L$ and widths $d_{1},\dots,d_{L}$ is any function $f:\mathbb{R}^{d}\to\mathbb{R}^{p}$ of the form $f(x)=x^{(L+1)}$ where
$\begin{cases}x^{(1)}=W^{(1)}x+b^{(1)}&\in\mathbb{R}^{d_{1}}\\[5.0pt] x^{(\ell+1)}=W^{(\ell+1)}\sigma(x^{(\ell)})+b^{(\ell+1)}&\in\mathbb{R}^{d_{\ell+1}}\qquad\ell=1,\dots,L\end{cases}$ (1)
for some weights $W^{(\ell)}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$ and biases (or offsets) $b^{(\ell)}\in\mathbb{R}^{d_{\ell}}$, and where the function $\sigma$ is assumed to apply on vectors component by component. We call a neural network shallow (or a one-hidden layer network) if $L=1$, deep if $L>1$. A neuron is a function of the form $\phi(z)=\sigma(\langle z,w\rangle+b)$. The vector $\sigma(x^{(\ell)})$ for $\ell=1,\dots,L$ is the $\ell$-th hidden layer of the network, hence $d_{\ell}$ is the number of neurons at the $\ell$-th hidden layer.
The input and output dimensions $d,p$ are prescribed by the problem. Given the problem, the user chooses an activation $\sigma$, a depth $L$ and widths $d_{1},\dots,d_{L}$. After this choice, the architecture is fixed, and the parameters $W^{(\ell)},b^{(\ell)}$ are obtained optimizing an empirical loss on the available data. Optimization is thus performed on the set of functions (1), with $(W^{(\ell)},b^{(\ell)})$ in $\mathbb{R}^{d_{\ell}\times d_{\ell-1}}\times\mathbb{R}^{d_{\ell}}$. The set of neural networks with a fixed architecture is parametric, albeit hence nonlinear and high-dimensional. Next, we will define structured nonparametric linear spaces on which optimization has a solution in the set of neural networks (1). This suggests that neural networks may be seen as parametric solutions of a nonparametric model, where the architecture itself can be obtained by training, rather than arbitrarily chosen by the practitioner. To define such a model, we need the notion of a reproducing kernel Banach space and vector measures.
Reproducing kernel Banach spaces
We will assume the following minimal definition, that readily generalizes reproducing kernel Hilbert spaces to a Banach setting. We refer to [30, 14] for a comprehensive overview.
Definition 2.2 (Vector valued reproducing kernel Banach space).
Let $\mathcal{X}$ be a set and $\mathcal{Y}$ a Banach space. A reproducing kernel Banach space (RKBS) $\mathcal{H}$ on $\mathcal{X}$ with values in $\mathcal{Y}$ is a Banach space of functions $f:\mathcal{X}\to\mathcal{Y}$ with continuous pointwise evaluation, that is, for all $x\in\mathcal{X}$ there is $C_{x}>0$ such that $\|f(x)\|_{\mathcal{Y}}\leq C_{x}\|f\|_{\mathcal{H}}$ for all $f\in\mathcal{H}$. In other words, the map $f\mapsto f(x)$ is in $B(\mathcal{H},\mathcal{Y})$ for all $x\in\mathcal{X}$.
The following proposition characterizes RKBS as Banach feature spaces, that is, spaces parametrized on Banach spaces.
Proposition 2.3.
Let $\mathcal{X}$ be a set and $\mathcal{Y}$ a Banach space. Consider the following statements.
  1. (a)
    A space $\mathcal{H}$ of functions $f:\mathcal{X}\to\mathcal{Y}$ is a RKBS.
  2. (b)
    There is a Banach space $\mathcal{F}$ and a map $\phi:\mathcal{X}\to B(\mathcal{F},\mathcal{Y})$ such that
    1. (i)
      =$\mathcal{H}=\{f_{\mu}:\mu\in\mathcal{F}\}$ where $f_{\mu}=\phi(\cdot)\mu$ ;
    2. (ii)
      $\|f\|_{\mathcal{H}}=\inf\{\|\mu\|_{\mathcal{F}}:\mu\in\mathcal{F},f=f_{\mu}\}$
  3. (c)
    There is a Banach space $\mathcal{F}$ and a map $\psi:\mathcal{X}\to B(\mathcal{Y}^{\prime},\mathcal{F}^{\prime})$ such that
    1. (i)
      =$\mathcal{H}=\{f_{\mu}:\mu\in\mathcal{F}\}$ where ${}_{\mathcal{Y}}\langle f_{\mu}(\cdot),y^{\prime}\rangle_{\mathcal{Y}^{\prime}}={}_{\mathcal{F}}\langle\mu,\psi(\cdot)y^{\prime}\rangle_{\mathcal{F}^{\prime}}$.
    2. (ii)
      $\|f\|_{\mathcal{H}}=\inf\{\|\mu\|_{\mathcal{F}}:\mu\in\mathcal{F},f=f_{\mu}\}$
Then (a) and (b) are equivalent and each one implies (c). Moreover, if $Y$ is reflexive (in particular Hilbert), then (a), (b) and (c) are all equivalent.
Proof.
To see that (a) implies (b), take $\mathcal{F}=\mathcal{H}$ and define
$\phi:\mathcal{X}\to B(\mathcal{F},\mathcal{Y})\qquad\phi(x)f=f(x)\ .$
Then (b)(i) and (b)(ii) of item (b) are clear. Let us prove that (b) implies (a). Clearly $\mathcal{H}$ is a linear space and $\|\cdot\|_{\mathcal{H}}$ is a norm. We then show that the normed space $\mathcal{H}$ is complete. The linear map $\mu\mapsto f_{\mu}$ has kernel $\mathcal{N}=\bigcap_{x\in\mathcal{X}}\ker\phi(x)$. Since $\phi(x)$ is bounded for all $x$, $\ker\phi(x)$ is closed, hence so is $\mathcal{N}$. Thus, $\mathcal{F}/\mathcal{N}$ is a Banach space [20, Theorem 1.41], by construction isomorphic to $\mathcal{H}$, which is therefore complete. Next, we show that point evaluations in $\mathcal{H}$ are continuous. For $f\in\mathcal{H}$, let $\mu\in\mathcal{F}$ such that $f=f_{\mu}$. Then
$\|f(x)\|_{\mathcal{Y}}=\|\phi(x)\mu\|_{\mathcal{Y}}\leq\|\phi(x)\|_{B(\mathcal{F},\mathcal{Y})}\|\mu\|_{\mathcal{F}}\ ,$
whence
$\|f(x)\|_{\mathcal{Y}}\leq\inf_{\mu\in\mathcal{F}:f=f_{\mu}}\|\phi(x)\|_{B(\mathcal{F},\mathcal{Y})}\|\mu\|_{\mathcal{F}}=\|\phi(x)\|_{B(\mathcal{F},\mathcal{Y})}\|f\|_{\mathcal{H}}\ .$
The implication from (b) to (c) follows easily considering $\psi(x)$ as the dual map of $\phi(x)$. Finally, note that (c)(i) in (c) defines $f_{\mu}$ as a function from $\mathcal{X}$ to $\mathcal{Y}^{\prime\prime}$. Hence, if $\mathcal{Y}$ is reflexive, it defines $f_{\mu}:\mathcal{X}\to\mathcal{Y}$. From here, following the proof of the implication from (b) to (a), one can prove that (c) implies (a). ∎
Vector Radon measures
In view of 2.3, RKBS can be constructed choosing a feature space parametrizing functions. Thinking of a neural network layer as an atomic integration, we will define RKBS parametrized by measures. As hidden layers have vectorial outputs, we need the concept of vector measure [10].
Definition 2.4 (Vector measures).
Let $\Theta$ be a (Hausdorff) locally compact, second countable topological space, and let $\mathcal{Y}$ be a Banach space. A vector (Radon) measure on $\Theta$ with values in $\mathcal{Y}$ is a map $\mu\colon\mathcal{B}(\Theta)\to\mathcal{Y}$ such that, for all $A\in\mathcal{B}(\Theta)$ and all $\{A_{i}\}$ partitions of $A$, [^1] A partition of a Borel set $A$ is a numerable family of Borel sets $\{A_{i}\}$ such that $A_{i}\cap A_{j}=\emptyset$ for all $i\neq j$ and $\bigcup_{i}A_{i}=A$.
$\mu(A)=\sum_{i}\mu(A_{i})\ ,$
where the sum converges absolutely in the $\|\cdot\|_{\mathcal{Y}}$ norm. We denote by $\mathcal{M}(\Theta,\mathcal{Y})$ the space of vector measures on $\Theta$ with values in $\mathcal{Y}$. If $\mu\in\mathcal{M}(\Theta,\mathcal{Y})$, for all $A\in\mathcal{B}(\Theta)$ we define
$|\mu|(A)=\sup_{\{A_{i}\}}\sum_{i}\|\mu(A_{i})\|_{\mathcal{Y}}\ ,$
where the supremum is taken over all partitions of $A$. Then, $|\mu|$ is a bounded positive measure on $\Theta$, which we call the total variation of $\mu$. The space $\mathcal{M}(\Theta,\mathcal{Y})$ is a Banach space with respect to the total variation norm
$\|\mu\|_{\operatorname{TV}}=|\mu|(\Theta)\ .$
Remark 2.5.
Fix a vector measure $\mu$. Since $\Theta$ is a second countable space, for all $y^{\prime}\in\mathcal{Y}^{\prime}$, ${}_{\mathcal{Y}}\langle\mu(\cdot),y^{\prime}\rangle_{\mathcal{Y}^{\prime}}$ is a scalar regular (Radon) measure, hence $\mu$ is a regular vector measure [21]. Since the elements of $\mathcal{M}(\Theta,\mathcal{Y}^{\prime})$ are regular measures, the space $\mathcal{M}(\Theta,\mathcal{Y}^{\prime})$ can be identified with the dual of $\mathcal{C}_{0}(\Theta,\mathcal{Y})$, the space of continuous functions from $\Theta$ to $\mathcal{Y}$ that vanish at infinity, by (a generalization of) the Riesz representation theorem. In particular, $\mathcal{M}(\Theta,\mathcal{Y}^{\prime})$ can be endowed with the product $*{}^{*}$ topology, with respect to which the closed balls are compact. This result was proved in [23] for compact $\Theta$ (see also [21]), and then extended in [8] to locally compact $\Theta$ in the case where $\mathcal{Y}$ is Hilbert. The proof in [8] holds true with the obvious modifications if $\mathcal{Y}$ is Banach.
3. Neural RKBS
In this section, we introduce particular classes of RKBS, incorporating general properties inspired by neural networks architectures. In particular, we construct deep and neural RKBS, and show that neural networks as intended in 2.1 are elements of neural RKBS. In this way, we provide neural networks with a Banach space functional structure, viewing them as finite representations of an underlying nonparametric model.
Integral RKBS
Let $\Theta,\Xi$ be two locally compact, second countable topological spaces. Let $\mathcal{X}$ be a set, and let $\mathcal{Y}$ be a RKBS on $\Xi$. Let
$\rho:\mathcal{X}\times\Theta\to\mathbb{R}$
be such that $\rho(x,\cdot)\in\mathcal{C}_{0}(\Theta)$ for all $x\in\mathcal{X}$. For $\mu\in\mathcal{M}(\Theta,\mathcal{Y})$, let
$f_{\mu}:\mathcal{X}\to\mathcal{Y}\qquad f_{\mu}(x)=\phi(x)\mu=\int_{\Theta}\rho(x,\theta){\rm d}\mu(\theta)$ (2)
and
=$\mathcal{H}=\{f_{\mu}:\mathcal{X}\to\mathcal{Y}:\mu\in\mathcal{M}(\Theta,\mathcal{Y})\}$
with
$\|f\|_{\mathcal{H}}=\inf_{\mu\in\mathcal{M}(\Theta,\mathcal{Y})}\{\|\mu\|_{\operatorname{TV}}:f=f_{\mu}\}\ .$ (3)
The space $\mathcal{H}$ thus defined is a (vector valued) RKBS, which we call an integral RKBS. Moreover, we call $\Theta$ and $\Xi$ the parameter spaces of $\mathcal{H}$, and $\rho$ its basis functions.
Since $\mathcal{Y}$ is a RKBS, for every $\xi\in\Xi$ there is $\delta_{\xi}\in\mathcal{Y}^{\prime}$ such that $y(\xi)={}_{\mathcal{Y}}\langle y,\delta_{\xi}\rangle_{\mathcal{Y}^{\prime}}$ for all $y\in\mathcal{Y}$. Thus, defining the (scalar) measure
$\mu(E\mid\xi)={}_{\mathcal{Y}}\langle\mu(E),\delta_{\xi}\rangle_{\mathcal{Y}^{\prime}}\qquad E\in\mathcal{B}(\Theta)\qquad\xi\in\Xi\ ,$
for all $x\in\mathcal{X}$ we have
$f_{\mu}(x)(\xi)=\int_{\Theta}\rho(x,\theta){\rm d}\mu(\theta\mid\xi)\ .$
Deep RKBS
Let $L$ be a positive integer. Take a set $\mathcal{X}_{0}$ and Banach spaces $\mathcal{X}_{1},\dots,\mathcal{X}_{L+1}$. For $\ell=0,\dots,L$, take RKBS $\mathcal{H}_{\ell}$ on $\mathcal{X}_{\ell}$ with values in $\mathcal{X}_{\ell+1}$. The direct sum
$\mathcal{H}=\mathcal{H}_{0}\oplus\cdots\oplus\mathcal{H}_{L}$
is a Banach space with respect to the norm
$\|f\|_{\mathcal{H}}=\|f_{0}\|_{\mathcal{H}_{0}}+\cdots+\|f_{L}\|_{\mathcal{H}_{L}}\qquad f=f_{0}\oplus\cdots\oplus f_{L}\ .$
To every $f=f_{0}\oplus\cdots\oplus f_{L}\in\mathcal{H}$ we assign the function $f^{\text{deep}}:\mathcal{X}_{0}\to\mathcal{X}_{L+1}$ defined by
$f^{\text{deep}}=f_{L}\circ\cdots\circ f_{0}\ .$
The linear space $\mathcal{H}$ parametrizes the nonlinear space
$\mathcal{H}^{\text{deep}}=\{f^{\text{deep}}:\mathcal{X}_{0}\to\mathcal{X}_{L+1}:f\in\mathcal{H}\}\ .$
With slight abuse of language, we call the non-linear space $\mathcal{H}^{\text{deep}}$ a RKBS of depth $L$ parametrized by the RKBS $\mathcal{H}=\mathcal{H}_{0}\oplus\cdots\oplus\mathcal{H}_{L}$. If $L>1$, we refer to $\mathcal{H}^{\text{deep}}$ as a deep RKBSp.
Moreover, we call $\mathcal{X}_{\ell}$ the layer spaces of $\mathcal{H}$. In particular, $\mathcal{X}=\mathcal{X}_{0}$ is the input space, $\mathcal{Y}=\mathcal{X}_{L+1}$ is the output space, and $\mathcal{X}_{\ell}$, $\ell=1,\dots,L$, are the hidden layer spaces.
Deep integral RKBS
Let $L$ be a positive integer. If the RKBS $\mathcal{H}_{\ell}$ are integral,
=$\mathcal{H}_{\ell}=\{f_{\mu_{\ell}}:\mathcal{X}_{\ell}\to\mathcal{X}_{\ell+1}:\mu_{\ell}\in\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})\}\ ,$
with basis functions
$\rho_{\ell}:\mathcal{X}_{\ell}\times\Theta_{\ell}\to\mathbb{R}\ ,$
the corresponding nonlinear space
$\mathcal{H}^{\text{deep}}=\{f^{\text{deep}}:\mathcal{X}_{0}\to\mathcal{X}_{L+1}:f^{\text{deep}}=f_{L}\circ\cdots\circ f_{0}\ \ ,f_{\ell}=f_{\mu_{\ell}}\in\mathcal{H}_{\ell}\ ,\ell=0,\ldots,L\}$
is called an integral RKBS of depth $L$ parametrized by $\mathcal{H}=\mathcal{H}_{0}\oplus\cdots\oplus\mathcal{H}_{L}$ and, if $L>1$, a deep integral RKBS.
A function $f:\mathcal{X}_{0}\to\mathcal{X}_{L+1}\in\mathcal{H}^{\text{deep}}$ has the form $f(x)=x^{(L+1)}$ where
$\begin{cases}x^{(0)}=x&\in\mathcal{X}_{0}\\[5.0pt] x^{(\ell+1)}=\displaystyle{\int_{\Theta_{\ell}}\rho_{\ell}(x^{(\ell)},\theta_{\ell}){\rm d}\mu_{\ell}(\theta_{\ell})}&\in\mathcal{X}_{\ell+1}\qquad\ell=0,\dots,L\ .\end{cases}$
Neural RKBS
We are going to define a particular instance of deep integral RKBS modeling on neural networks. In neural RKBS, the basis functions are defined by activation functions. First, we fix the parameter spaces $\Theta_{\ell}$ and the layer spaces $\mathcal{X}_{\ell}$ as follows:
$\Theta_{0}=\{0,\dots,d\}\qquad$ $\mathcal{X}_{0}=\ell^{2}(\{1,\dots,d\})=\mathbb{R}^{d}\qquad$
$\Theta_{\ell}=\mathbb{N}\qquad$ $\mathcal{X}_{\ell}=\ell^{2}(\mathbb{N})\qquad$ $\ell=1,\dots,L$
$\Theta_{L+1}=\{1,\dots,p\}\qquad$ $\mathcal{X}_{L+1}=\ell^{2}(\{1,\dots,p\})=\mathbb{R}^{p}\qquad$ $\ell=L+1\ .$
Note that $\mathcal{X}_{0}$ can be thought of as a function space on $\Theta_{0}$ by putting $x(0)=1$ for all $x\in\mathcal{X}_{0}$.
Next, let $\sigma:\mathbb{R}\to\mathbb{R}$ be a Lipschitz activation function such that $\sigma(0)=0$. This implies that
$\sigma(x)\in\ell^{2}(\mathbb{N})\qquad\text{for all }x\in\ell^{2}(\mathbb{N})\ ,$ (4)
where $\sigma$ applies on sequences component by component. Indeed, if $C_{\sigma}$ denotes the Lipschitz constant of $\sigma$, we have
$|\sigma(x_{n})|$ $=|\sigma(x_{n})-\sigma(0)|\leq C_{\sigma}|x_{n}-0|=C_{\sigma}|x_{n}|.$
In particular, this implies that
$\lim_{n\to\infty}\sigma(x_{n})=0\qquad\text{for all }x\in\ell^{2}(\mathbb{N})\ .$ (5)
Assuming $\sigma(0)=0$ is not restrictive. In fact, if $\sigma(0)\neq 0$, the choice of the activation function $\sigma^{\prime}=\sigma-\sigma(0)$ simply implies a scaling of the offsets. Then, we set
$\rho_{0}(x,n)=\begin{cases}1&n=0\\ x_{n}&n=1,\ldots,d\end{cases}\qquad x\in\mathbb{R}^{d}$
$\rho_{\ell}(x,n)=\begin{cases}1&n=0\\ \sigma(x_{n-1})&n\geq 1\end{cases}\quad x\in\ell^{2}(\mathbb{N})$ $\qquad\ell=1,\dots,L\ .$
We call the resulting (deep) integral RKBS $\mathcal{H}^{\text{deep}}$ a neural RKBS of depth $L$ and, if $L>1$, a deep neural RKBS.
Remark 3.1.
If $\sigma$ is Lipschitz with Lipschitz constant $C_{\sigma}$, then, for all $\ell=1,\dots,L$, $x,x^{\prime}\in\ell^{2}(\mathbb{N})$ and $n\in\mathbb{N}$,
$|\rho_{\ell}(x,n)-\rho_{\ell}(x^{\prime},n)|\leq|\sigma(x_{n-1})-\sigma(x^{\prime}_{n-1})|\leq C_{\sigma}|x_{n-1}-x^{\prime}_{n-1}|.$
General form of neural RKBS functions
An element $f^{\text{deep}}$ of the neural RKBS $\mathcal{H}^{\text{deep}}$ of depth $L$ is a composition of $L+1$ integral functions $f_{0}$,…,$f_{L}$,
[…]
where each $f_{\ell}$ is defined via a measure $\mu_{\ell}$ by $f_{\ell}=f_{\mu_{\ell}}$. From layer 00 to layer $1$, we have
$\mu_{0}\in\mathcal{M}(\{0,\dots,d\},\ell^{2}(\mathbb{N}))\ ,$
so that
$\quad\mu_{0}=\sum_{m=0}^{d}w^{(1)}_{m}\delta_{m}$
for a $d+1$-family of vectors $w^{(1)}_{0},\ldots w^{(1)}_{d}\in\ell^{2}(\mathbb{N})$. Let $b^{(1)}=w^{(1)}_{0}\in\ell^{2}(\mathbb{N})$ and define the bounded operator
$W^{(1)}\in B(\mathbb{R}^{d},\ell^{2}(\mathbb{N}))\qquad W^{(1)}x=\sum_{m=1}^{d}w^{(1)}_{m}x_{m}\ ,$
Then, for $x\in\mathbb{R}^{d}$, the function $f_{0}:\mathbb{R}^{d}\to\ell^{2}(\mathbb{N})$ is
$f_{0}(x)=\sum_{m=0}^{d}w^{(1)}_{m}\rho_{0}(x,m)=w^{(1)}_{0}+\sum_{m=1}^{d}w^{(1)}_{m}x_{m}=W^{(1)}x+b^{(1)}\in\ell^{2}(\mathbb{N})\ ,$
and the scalar components of $f_{0}$ are
$f_{0}(x)_{n}=\langle x,w^{(1)}_{n\cdot}\rangle_{\mathbb{R}^{d}}+b^{(1)}_{n}\qquad n\in\mathbb{N}\ ,$
where $w^{(1)}_{n\cdot}\in\mathbb{R}^{d}$ with $w^{(1)}_{nm}=(w^{(1)}_{m})_{n}$.
For $\ell=1,\dots,L-1$, from layer $\ell$ to layer $\ell+1$ we have
$\mu_{\ell}\in\mathcal{M}(\mathbb{N},\ell^{2}(\mathbb{N}))\ ,$
so that
$\mu_{\ell}=\sum_{m=0}^{\infty}w^{(\ell+1)}_{m}\delta_{m}\ ,$
for a countable family of vectors $w^{(\ell+1)}_{0},\ldots,w^{(\ell+1)}_{m},\ldots\in\ell^{2}(\mathbb{N})$ such that
$\|\mu_{\ell}\|_{\text{TV}}=\sum_{m=0}^{\infty}\|w^{(\ell+1)}_{m}\|_{\ell^{2}(\mathbb{N})}<+\infty\ .$ (6)
As before, set $b^{(\ell+1)}=w^{(\ell+1)}_{0}\in\ell^{2}(\mathbb{N})$ and
$W^{(\ell+1)}\in B(\ell^{2}(\mathbb{N}),\ell^{2}(\mathbb{N}))\qquad W^{(\ell+1)}x=\sum_{m=1}^{+\infty}w^{(\ell+1)}_{m}x_{m-1}\ ,$
where the series converges absolutely in $\ell_{2}(\mathbb{N})$ due to (6). Hence, for $x\in\ell_{2}(\mathbb{N})$,
$f_{\ell}(x)$ $=\sum_{m=0}^{+\infty}w^{(\ell+1)}_{m}\rho_{\ell}(x,m)=w^{(\ell+1)}_{0}+\sum_{m=1}^{+\infty}w^{(\ell+1)}_{m}\sigma(x_{m-1})$
$=W^{(\ell+1)}\left(\sigma(x)\right)+b^{(\ell+1)}\in\ell^{2}(\mathbb{N})\ ,$
where $\sigma(x)\in\ell^{2}(\mathbb{N})$ by (4). The component of $f_{\ell}(x)$ are
$f_{\ell}(x)_{n}=\langle\sigma(x),w^{(\ell+1)}_{n\cdot}\rangle_{\ell^{2}}+b^{(\ell+1)}_{n}\qquad n\in\mathbb{N}\ ,$
where $w^{(\ell+1)}_{n\cdot}\in\ell^{2}(\mathbb{N})$ and $w^{(\ell+1)}_{nm}=(w^{(\ell+1)}_{m})_{n}$.
Finally, from layer $L$ to layer $L+1$, we have
$\mu_{L}\in\mathcal{M}(\mathbb{N},\mathbb{R}^{p})\ ,$
so that
$\mu_{L}=\sum_{m=0}^{\infty}w^{(L+1)}_{m}\delta_{m}\ ,$
for a countable family of vectors $w^{(L+1)}_{0},\ldots,w^{(L+1)}_{m},\ldots\in\mathbb{R}^{p}$ such that
$\sum_{m=0}^{\infty}\|w^{(\ell+1)}_{m}\|_{\mathbb{R}^{p}}<+\infty\ .$ (7)
As before, set $b^{(L+1)}=w^{(L+1)}_{0}\in\ell^{2}(\mathbb{N})$ and
$W^{(L+1)}\in B(\ell^{2}(\mathbb{N}),\mathbb{R}^{p})\qquad W^{(L+1)}x=\sum_{m=1}^{+\infty}w^{(L+1)}_{m}x_{m-1}\ ,$
where the series converges absolutely in $\mathbb{R}^{p}$ due to (7). Hence, for $x\in\ell_{2}(\mathbb{N})$,
$f_{L}(x)$ $=\sum_{m=0}^{+\infty}w^{(L+1)}_{m}\rho_{L}(x,m)=w^{(L+1)}_{0}+\sum_{m=1}^{+\infty}w^{(L+1)}_{m}\sigma(x_{m-1})$
$=W^{(L+1)}\left(\sigma(x)\right)+b^{(L+1)}\in\ell^{2}(\mathbb{N})\ ,$
where $\sigma(x)\in\ell^{2}(\mathbb{N})$ by (4). The component of $f_{L}(x)$ are
$f_{L}(x)_{n}=\langle\sigma(x),w^{(L+1)}_{n\cdot}\rangle_{\ell^{2}}+b^{(L+1)}_{n}\qquad n=1,\ldots,p\ ,$ (8)
where $w^{(L+1)}_{n\cdot}\in\ell^{2}(\mathbb{N})$ and $w^{(L+1)}_{nm}=(w^{(L+1)}_{m})_{n}$.
By iteration $x^{(\ell+1)}=f_{\ell}(x^{(\ell)})$, we obtain
$\begin{cases}x^{(0)}=x&\in\mathbb{R}^{d}\\[5.0pt] x^{(1)}=W^{(1)}x^{(0)}+b^{(1)}&\in\ell^{2}(\mathbb{N})\\[5.0pt] x^{(\ell+1)}=W^{(\ell+1)}\left(\sigma(x^{(\ell)})\right)+b^{(\ell+1)}&\in\ell^{2}(\mathbb{N})\\[5.0pt] x^{(L+1)}=W^{(L+1)}\left(\sigma(x^{(L)})\right)+b^{(L+1)}&\in\mathbb{R}^{p}\ .\end{cases}$
We call the neural RKBS function $f$ an infinite-width neural network. Note that $\mathbb{N}$ parametrizes the width of the $L$ hidden layers $\ell=1,\dots,L$, while input and output layers $\ell=0$ and $\ell=L+1$ have fixed finite widths $d$ and $p$, respectively. Also note that infinite-width neural networks generalize finite-width neural networks, where the inner layers are generated by infinite-rank operators.
We can visualize the shallow ($L=1$), and the simplest non-shallow case ($L=2$), considering a non-iterative expression. For $L=1$, we have
$f(x)=W^{(2)}\left(\sigma(W^{(1)}x+b^{(1)})\right)+b^{(2)}\ .$
In the case of $L=2$ hidden layers, we can write
$f(x)=W^{(3)}\left(\sigma\left(W^{(2)}\left(\sigma(W^{(1)}x+b^{(1)})\right)+b^{(2)}\right)\right)+b^{(3)}\ .$
Comparing our construction to the one proposed in [18], we remark that our networks are completely nonparametric, in the sense that every hidden layer has infinte width. In contrast, the networks in [18] are compositions of vector-valued shallow networks hence they have intermediate linear layers of finite width, corresponding to the finite output dimension of each shallow network. Such an architecture can be reinterpreted as an infinite-width network with finite-rank constraints. While imposing such constraints may enforce in practice some sort of beneficial stability, how to set or bound the ranks is an arbitrary choice.
Finite form of neural RKBS functions
We now show that the neural functions defined in 2.1 are correspond to measures $\mu_{1},\ldots,\mu_{L}$ having finite support. Indeed, under this assumption, for each $\ell=1,\ldots,L$,
$\mu_{\ell}=b^{(\ell+1)}\delta_{0}+\sum_{k=1}^{d_{\ell}}w^{(\ell+1)}_{k}\delta_{m^{(\ell)}_{k}}$
for some family of location $m^{(\ell)}_{1},\ldots,m^{(\ell)}_{d_{\ell}}\in\mathbb{N}\setminus\{0\}$ and some family of vectors $w^{(\ell+1)}_{1},\ldots,w^{(\ell+1)}_{k}\in\ell_{2}(\mathbb{N})$ if $\ell<L$ ($w^{(L+1)}_{1},\ldots,w^{(L+1)}_{k}\in\mathbb{R}^{p}$ if $\ell=L$). We define $f_{\ell}$ starting from the last layer. Since the support of $\mu_{L}$ is $\{0,m^{{L}}_{1},\ldots,m^{(L)}_{d_{L}}\}$, by (8) $f_{L}$ depends only on the variables $m^{(L)}_{1},\ldots,m^{(L)}_{d_{L}}$, so that we can regard $f_{L}$ as a function from $\mathbb{R}^{d_{L}}$ to $\mathbb{R}^{p}$ given by
$f_{L}(x)=W^{(L+1)}\sigma(x)+b\ ,$
where $W^{(L+1)}$ is the $d_{L}\times d_{L+1}$ matrix (recall that $d_{L+1}=p$) with components
$W^{(L+1)}_{nk}=(w^{(L+1)}_{m_{k}^{(L)}})_{n-1}\qquad n=1,\ldots,p,\ k=1,\ldots,d_{L}\ .$
Since $f_{L}$ is defined on $\mathbb{R}^{d_{L}}$, regarded as finite dimensional subspace of $\ell_{2}(\mathbb{N})$, and denoted by $P:\ell_{2}(\mathbb{N})\to\mathbb{R}^{d_{L}}$ the corresponding projection
$Px=(x_{m^{(L)}_{1}},\ldots,x_{m^{(L)}_{d_{L}}})\ ,$
then for all $x\in\ell_{2}(\mathbb{N})$
$f_{L}(f_{L-1}(x))=f_{L}(f_{\mu_{L-1}}(x))=f_{L}(f_{P\mu_{L-1}}(x))\ .$
Then, without loss of generality, we can assume that the measure $\mu_{L-1}$ is in $\mathcal{M}(\mathbb{N},\mathbb{R}^{d_{L}})$ and it has a finite support. By iterating this procedure, we can assume that for all $\ell$
$\mu_{\ell}\in\mathcal{M}(\{0,\ldots,d_{\ell}\},\mathbb{R}^{d_{\ell+1}})$
for some family $d_{0},d_{1},\dots,d_{L},d_{L+1}\in\mathbb{N}$ with $d_{0}=d$ and $d_{L+1}=p$. This means that
$\mu_{\ell}=b^{(\ell+1)}\delta_{0}+\sum_{k=1}^{d_{\ell}}w^{(\ell+1)}_{k}\delta_{k}\ .$
Hence, for all $\ell=1,...,L+1$, let $W^{(\ell)}$ be the $n_{d_{\ell-1}}\times n_{d_{\ell}}$ matrix
$W^{(\ell)}_{nk}=(w^{(\ell)}_{k})_{n}\qquad n=1,\ldots,d_{\ell},\ k=1,\ldots,d_{\ell-1}\ .$
Then $f_{\ell}=f_{\mu_{\ell}}:\mathbb{R}^{d_{\ell-1}}\to\mathbb{R}^{d_{\ell}}$ is given by
$f_{\ell}(x)=\begin{cases}W^{(1)}x+b^{(1)}&\ell=0\\ W^{(\ell+1)}\sigma(x)+b^{(\ell+1)}&\ell>1\ ,\end{cases}$
$f^{\text{deep}}=f_{L}\circ\ldots\circ f_{0}$
is a neural deep function according to 2.1, and we can rewrite $f^{\text{deep}}$ as
$\begin{cases}x^{(1)}=W^{(1)}x+b^{(1)}&\in\mathbb{R}^{d_{1}}\\[5.0pt] x^{(\ell+1)}=W^{(\ell+1)}\sigma(x^{(\ell)})+b^{(\ell+1)}&\in\mathbb{R}^{d_{\ell+1}}\qquad\ell=1,\dots,L\ ,\end{cases}$
that is, $f^{\text{deep}}$ is a neural network of depth $L$ and (finite) widths $d_{1},\dots,d_{L}$.
4. Representer theorems
In this section, we state and prove representer theorems for deep and neural RKBS, showing that neural networks can be seen as parametric solutions of a nonparametric model. We start by briefly recalling a basic supervised learning setting.
Let $\mathcal{X}$ be a set and $\mathcal{Y}$ a Banach space, called input and output space, respectively. Consider a Banach space $\mathcal{H}$ of functions $f:\mathcal{X}\to\mathcal{Y}$, called hypothesis space, and a loss function $\mathcal{L}:\mathcal{Y}\times\mathcal{Y}\to[0,\infty)$. Given $N$ samples
$(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\qquad i=1,\dots,N\ ,$
we want to solve the regularized empirical risk minimization problem
$\inf_{f\in\mathcal{H}}\mathcal{R}(f)+\|f\|_{\mathcal{H}}\ ,$ (9)
where
$\mathcal{R}(f)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f(x_{i}),y_{i})\ ,$
is the empirical error associated to the loss function $\mathcal{L}$ and the training points $(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y},i=1,\dots,N$.
To prove a representer theorem when $\mathcal{H}$ is a deep integral RKBS, we will need the following two fundamental assumptions on the basis functions $\rho_{\ell}$. We will see that both assumptions are easily satisfied by neural RKBS.
Assumption 1 (Pointwise form).
There is $\widetilde{\rho}_{\ell}:\mathbb{R}\times\Theta_{\ell}\to\mathbb{R}$ such that, for all $x\in\mathcal{X}_{\ell}$ and $\theta\in\Theta_{\ell}$,
$\rho_{\ell}(x,\theta)=\widetilde{\rho}_{\ell}(x(\theta),\theta)\ .$
Assumption 2 (Lipschitz condition).
There are $C_{\ell}>0$, $g_{\ell}\in\mathcal{C}_{b}(\Theta_{\ell},\mathcal{X}_{\ell})$ and $\beta_{\ell}\in\mathcal{C}_{0}(\Theta_{\ell})$ such that, for all $x,x^{\prime}\in\mathcal{X}_{\ell}$ and $\theta\in\Theta_{\ell}$,
$|\rho_{\ell}(x,\theta)-\rho_{\ell}(x^{\prime},\theta)|\leq C_{\ell}|\langle x-x^{\prime},g_{\ell}(\theta)\rangle_{\mathcal{X}_{\ell}}||\beta_{\ell}(\theta)|\ .$
For neural RKBS with activation function $\sigma$, 1 is satisfied with $\widetilde{\rho}_{0}(t,n)=t$ and $\widetilde{\rho}_{\ell}(t,n)=\sigma(t)$ for all $\ell=1,\dots,L$. Furthermore, by Remark 3.1, 2 is satisfied with $C_{\ell}=C_{\sigma}$ and $g_{\ell}(n)=\delta_{n-1}$ for all $\ell=1,\dots,L$.
Theorem 4.1 (Representer theorem for deep integral RKBS).
Let $\mathcal{H}$ be an integral RKBS of depth $L$ on $\mathcal{X}$ with values in $\mathcal{Y}$. Assume that $\mathcal{X}_{\ell+1}$, $\ell=0,\dots,L$, is Hilbert, $\mathcal{Y}=\mathcal{X}_{L+1}$ has finite dimension $d_{L+1}$, and $\mathcal{L}$ is continuous in the first entry. Moreover, assume that $\rho_{\ell}$ satisfies 1 for all $\ell=0,\dots,L$ and 2 for all $\ell=1,\dots,L$. Then, there are $d_{1},\dots,d_{L}<\infty$, $\theta^{(\ell)}\in\Theta_{\ell}^{d_{\ell}}$ and $w^{(\ell)}\in\mathcal{X}_{\ell}^{d_{\ell-1}}$ such that $f(x)=x^{(L+1)}$ defined by
$\begin{cases}x^{(0)}=x\\[5.0pt] x^{(\ell+1)}=\sum_{k=1}^{d_{\ell}}w^{(\ell+1)}_{k}\rho_{\ell}(x^{(\ell)},\theta^{(\ell)}_{k})\qquad\ell=0,\dots,L\end{cases}$
is a solution to problem (9). Moreover, for every $\ell=1,\dots,L$, $d_{\ell}\leq Nd_{\ell+1}$, and
$\|f\|_{\mathcal{H}}\leq\sum_{\ell=0}^{L}\sum_{k=1}^{d_{\ell}}\|w^{(\ell+1)}_{k}\|_{\mathcal{X}_{\ell+1}}\ .$
Proof.
Recall that in (9) the evaluation of an element $f=f_{0}\otimes f_{L}\in\mathcal{H}$ at point $x\in\mathcal{X}$ is given by $f(x_{i})=f^{\text{deep}}(x)$ where $f^{\text{deep}}$ is the composition of $f_{0},\ldots,f_{L}$. Furthermore, by construction each $f_{\ell}\in\mathcal{H}_{\ell}$ is parameterised by some measure $\mu_{\ell}\in\mathcal{M}_{\ell}=\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})$ according to (2). Taking into account account (3), the minimization problem (9) is equivalent to
$\inf_{\mu\in\mathcal{M}}\mathcal{R}(f_{\mu}^{\text{deep}})+\sum_{\ell=0}^{L}\|\mu_{\ell}\|_{\operatorname{TV}}=:\inf_{\mu\in\mathcal{M}}\mathcal{S}(\mu),$ (10)
where if $\mu=\mu_{0}\otimes\mu_{L}$, $f_{\mu}^{\text{deep}}$ is the composition of $f_{\mu_{0}},\ldots,f_{\mu_{L}}$.
By a classical argument, we first show that Problem (10) admits a solution. Pick a fixed $\nu\in\mathcal{M}$, and let $R=\mathcal{S}(\nu)$. Then Problem (10) is equivalent to
$\inf_{\mu\in\prod_{\ell=0}^{L}B_{\mathcal{M}_{\ell}}(R)}\mathcal{S}(\mu)\ .$ (11)
Indeed, take $\mu$ outside $\prod_{\ell=0}^{L}B_{\mathcal{M}_{\ell}}(R)$, then for some $\ell=0,\ldots,L$, $\|\mu_{\ell}\|_{\text{TV}}>R$, so that
$\mathcal{S}(\mu)=\mathcal{R}(f_{\mu})+\sum_{\ell=0}^{L}\|\mu_{\ell}\|_{\operatorname{TV}}\geq\mathcal{R}(f_{\mu})+\mathcal{S}(\nu)\geq\mathcal{S}(\nu)\ ,$
which proves the equivalence. We now prove the existence of a minimizer of (11). By A.7, for all $x\in\mathcal{X}$ the map
$(\mu_{0},\dots,\mu_{L})\mapsto f_{\mu_{L}}\circ\cdots\circ f_{\mu_{0}}(x)$
is jointly continuous from $\prod_{\ell=0}^{L}B_{\mathcal{M}_{\ell}}(R)$, endowed with product topology induced by the weak-* topology of each $B_{\mathcal{M}_{\ell}}(R)$ to $\mathcal{Y}$. Since the $\mu_{\ell}\to\|\mu\|_{\operatorname{TV}}$ is weak-* continuous, the map $\mu\mapsto\mathcal{S}(\mu)$ is also continuous. Moreover, thanks to the Banach–Alaoglu theorem, the product $\prod_{\ell=0}^{L}B_{\mathcal{M}_{\ell}}(R)$ is $*{}^{*}$ compact. Hence, by the extreme value theorem, the problem (11) has at least a minimizer.
Let $\mu^{*}$ be any such solution. Denote $x_{i}^{(0)}=x_{i}$ and $x_{i}^{(\ell+1)}=f_{\mu^{*}_{\ell}}(x_{i}^{(\ell)})$ for $\ell=0,\dots,L$. Then, in view of A.4, a solution to (11) can be found by solving the following interpolation problems for all $\ell=0,\dots,L$:
$\inf_{\mu_{\ell}\in\mathcal{M}_{\ell}}\|\mu_{\ell}\|_{\operatorname{TV}}\quad\text{subject to}\quad f_{\mu_{\ell}}(x_{i}^{(\ell)})=x_{i}^{(\ell+1)}\quad i=1,\dots,N\ .$ (12)
Let us start from $\ell=L$. We want to apply A.2. To this end, let $\mathcal{U}=\mathcal{M}_{L}$ endowed with the $*{}^{*}$ topology. We define $\mathcal{A}:\mathcal{U}\to\mathbb{R}^{N\times d_{L+1}}$ by
$\mathcal{A}\mu=[f_{\mu_{L}}(x_{i})]_{i=1,\dots,N}\ .$
Then $\mathcal{A}$ is a surjective continuous linear operator from $\mathcal{U}$ onto $H=\operatorname{Ran}\mathcal{A}$, with $\dim(H)\leq Nd_{L+1}$. Moreover, the norm $G=\|\cdot\|_{\operatorname{TV}}$ is coercive on $\mathcal{U}$. Indeed, by the Banach–Alaoglu theorem, the balls $B_{\mathcal{M}_{L}}(r)$ are $*{}^{*}$ compact for every $r>0$. We define $F:H\to[0,\infty]$ by
$F(h)=\begin{cases}0&h_{i}=y_{i}\text{ for all $i=1,\dots,N$}\\ \infty&h_{i}\neq y_{i}\text{ for some $i=1,\dots,N$}\ .\end{cases}$
The function $F$ is the indicator function associeated to the singleton $\{(y_{0},\ldots,y_{L})\}$, so that it is convex, coercive and lower semi-continuous. Therefore, we can apply A.2, whence we obtain that (12) has a solution of the form
$\widetilde{\mu}_{L}=\sum_{k=1}^{d_{L}}c_{k}u_{k}$
for some $d_{L}\leq Nd_{L+1}$, $c_{k}>0$ and $u_{k}\in\operatorname{Ext}(B_{\mathcal{M}_{L}}(1))$. By A.3, for each $k$ there are $y_{k}\in\operatorname{Ext}(B_{\mathcal{Y}}(1))$ and $\theta^{(L)}_{k}\in\Theta_{L}$ such that
$u_{k}=y_{k}\cdot\delta_{\theta^{(L)}_{k}}\ .$
Thus, defining $w^{(L+1)}_{k}=c_{k}y_{k}$, we get
$\widetilde{\mu}_{L}=\sum_{k=1}^{d_{L}}w^{(L+1)}_{k}\delta_{\theta^{(L)}_{k}}\ .$
We now come to $\ell=L-1$. Thanks to 1, we can regard $f_{\widetilde{\mu}_{L}}$ as defined on
$\widetilde{\mathcal{X}}_{L}=\mathbb{R}^{\operatorname{{supp}}(\widetilde{\mu}_{L})}\ ,$
and restrict (12) to
$\inf_{\mu_{L-1}\in\mathcal{M}(\Theta_{L-1},\widetilde{\mathcal{X}}_{L})}\|\mu_{L-1}\|_{\operatorname{TV}}\quad\text{subject to}\quad f_{\mu_{L-1}}(x_{i}^{(L-1)})={x_{i}^{(L)}}_{|\operatorname{{supp}}(\widetilde{\mu}_{L})}\quad i=1,\dots,N\ .$
We use again A.2, together with A.3, as in the previous step ($\ell=L$). In particular, this time we have $\dim(H)\leq Nd_{L}$. Thus, we find a solution of the form
$\widetilde{\mu}_{L-1}=\sum_{k=1}^{d_{L-1}}w^{(L)}_{k}\delta_{\theta^{(L-1)}_{k}}\ ,$
for some $d_{L-1}\leq Nd_{L}$.
Iterating the argument for $\ell=L-2,\dots,0$, the main claim follows. The bound on $d_{\ell}$ is also clear by iteration. For the bound on $\|f\|_{\mathcal{H}}$, by definition we have
$\|f\|_{\mathcal{H}}=\sum_{\ell=0}^{L}\|f_{\ell}\|_{\mathcal{H}_{\ell}}=\sum_{\ell=0}^{L}\|f_{\widetilde{\mu}_{\ell}}\|_{\mathcal{H}_{\ell}}\leq\sum_{\ell=0}^{L}\|\widetilde{\mu}_{\ell}\|_{\operatorname{TV}}\ ,$
where, once again by A.2,
$\|\widetilde{\mu}_{\ell}\|_{\operatorname{TV}}=\sum_{k=1}^{d_{\ell}}c_{k}\ .$
But since $y_{k}\in\operatorname{Ext}(B_{\mathcal{X}_{\ell+1}}(1))$, we have $\|y_{k}\|_{\mathcal{X}_{\ell+1}}=1$, whence
$\sum_{k=1}^{d_{\ell}}c_{k}=\sum_{k=1}^{d_{\ell}}\|c_{k}y_{k}\|_{\mathcal{X}_{\ell+1}}=\sum_{k=1}^{d_{\ell}}\|w^{(\ell+1)}_{k}\|_{\mathcal{X}_{\ell+1}}\ ,$
which concludes the proof. ∎
In the special case of a neural RKBS, 4.1 implies the following result.
Corollary 4.2 (Representer theorem for neural RKBS).
Let $\mathcal{H}$ be a neural RKBS of depth $L$ on $\mathbb{R}^{d}$ with valued in $\mathbb{R}^{d_{L+1}}$. Assume that $\mathcal{L}$ is continuous in the first entry. Then, there are $d_{1},\dots,d_{L}<\infty$, $W^{(\ell)}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$ and $b^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ such that $f(x)=x^{(L+1)}$ defined by
$\begin{cases}x^{(1)}=W^{(1)}x\\[5.0pt] x^{(\ell+1)}=W^{(\ell+1)}\sigma(x^{(\ell)}-b^{(\ell)})\qquad\ell=1,\dots,L\end{cases}$
is a solution to problem (9). Moreover, for every $\ell=1,\dots,L$, $d_{\ell}\leq Nd_{\ell+1}$, and
$\|f\|_{\mathcal{H}}\leq\sum_{\ell=0}^{L}\sum_{k=1}^{d_{\ell}}\biggl{(}\sum_{j=1}^{d_{\ell+1}}|W^{(\ell+1)}_{jk}|^{2}\biggr{)}^{1/2}\ .$
Our result shows that deep neural networks are solutions to regularized empirical risk minimization over neural RKBS, thus justifying neural RKBS as hypothesis spaces for deep learning. Moreover, it provides an upper bound on the network width depending on sample size and input/output dimensions. Finally, it shows that the regularization norm is controlled by the $\ell^{1}$ norm of the $\ell^{2}$ norms of the weights of the network. Finer characterizations of the Banach structure can be obtained using the specific form of the activation function and the functional properties that this induces (see [18] for the ReLU).
5. Conclusion
Studying function spaces defined by neural networks provides a natural way to understand their properties. Recently reproducing kernel Banach spaces have emerged has a useful concept to study shallow networks.
In this paper, we take a step towards more complex architectures considering deep networks. We allow for a wide class of activation functions and remove unnecessary low rank constraints. Our main contribution is defining a class of neural RKBS obtained composing vector valued RKBS and deriving corresponding representer theorems borrowing ideas from [18].
Future developments include considering more structured architectures, for example convolutional networks, as well as investigating the statistical and computational properties of neural RKBS
Appendix A Sparse solutions to finite-dimensional variational problems
The key ingredient to establish our representer theorem is given by a powerful variational result proved in [7]. This result deals with general minimization problems with finite-dimensional constraints and seminorm penalization. It states that such problems admit sparse solutions, namely finite linear combinations of extremal points of the seminorm unit ball. We report the formal statement below, after recalling the definition of extremal point.
Definition A.1 (Extremal point).
Let $Q$ be a convex subset of a locally convex space. A point $q\in Q$ is called extremal if $Q\setminus\{q\}$ is convex, that is, there do not exist $p,r\in Q$, $p\neq r$, such that $q=tp+(1-t)r$ for some $t\in(0,1)$. We denote the set of extremal points of $Q$ by $\operatorname{Ext}(Q)$.
Theorem A.2 ([7, Theorem 3.3]).
Consider the problem
$\operatorname*{arg\,min}_{u\in U}F(\mathcal{A}u)+G(u)\ ,$ (13)
where $U$ is a locally convex topological vector space, $\mathcal{A}:U\to H$ is a continuous, surjective linear map with values in a finite-dimensional Hilbert space $H$, $F:H\to(-\infty,+\infty]$ is proper, convex, coercive and lower semi-continuous, and $G:U\to[0,+\infty)$ is a coercive and lower semi-continuous norm. Let $B_{U}(1)$ denote the unit ball $\{u\in U:G(u)\leq 1\}$. Then (13) has solutions of the form
$\sum_{k=1}^{K}c_{k}u_{k}$
with $u_{k}\in\operatorname{Ext}(B_{U}(1))$, $c_{k}>0$, $K\leq\dim H$, and $\sum_{k=1}^{K}c_{k}=G(u)$.
A.2 is a simplified version of [7, Theorem 3.3], where $G$ is only assumed to be a seminorm.
In view of A.2, in order to determine the form of sparse solutions we need to characterize the set of extremal point of the unit ball in $U$. In our construction of integral RKBS, $U$ is the space of vector measures with total variation norm. The following result provides the desired characterization for our case. It appears in [29] assuming that $\Theta$ is compact, but the proof works analogously when $\Theta$ is locally compact. We report the proof for the reader’s convenience.
Lemma A.3 ([29, Theorem 2]).
Let $\Theta$ be a locally compact, second countable topological space, and let $\mathcal{Y}$ be a Banach space. Then
$\operatorname{Ext}(B_{\mathcal{M}(\Theta,\mathcal{Y})}(1))=\{y\cdot\delta_{\theta}:y\in\operatorname{Ext}(B_{\mathcal{Y}}(1)),\theta\in\Theta\}.$
Proof.
Let us denote $E=\{y\cdot\delta_{\theta}:y\in\operatorname{Ext}(B_{\mathcal{Y}}(1)),\theta\in\Theta\}$. We start showing that $\operatorname{Ext}(B_{\mathcal{M}}(1))\subseteq E$. Suppose that $\mu\in\operatorname{Ext}(B_{\mathcal{M}}(1))$ but $\mu\neq y\cdot\delta_{\theta}$ for any $y\in\operatorname{Ext}(B_{\mathcal{Y}}(1))$ and $\theta\in\Theta$. Then $|\mu|\neq\delta_{\theta}$ for any $\theta\in\Theta$, and there is $A\in\mathcal{B}(\Theta)$ such that $10<|\mu|(A)<$ 1. Denote by $\chi_{A}$ the indicator function on $A$. Then, setting $t=|\mu|(A)$, $\mu_{1}=\mu\chi_{A}/t$ and $\mu_{2}=\mu\chi_{\Theta\setminus A}/(1-t)$, we can write $\mu$ as a convex combination
$\mu=t\mu_{1}+(1-t)\mu_{2}\ .$ (14)
Since $t\in(0,1)$ and $\mu_{1},\mu_{2}\in B_{\mathcal{M}}(1)$, we get that $\mu\notin\operatorname{Ext}(B_{\mathcal{M}}(1))$, leading to a contradiction. We now show the converse inclusion $E\subseteq\operatorname{Ext}(B_{\mathcal{M}}(1))$. Let $\mu=y\cdot\delta_{\theta}$ for some $y\in\operatorname{Ext}(B_{\mathcal{Y}}(1))$ and $\theta\in\Theta$. Suppose there are $t\in(0,1)$ and $\mu_{1},\mu_{2}\in B_{\mathcal{M}}(1)$ such that (14). We want to show that necessarily $\mu_{1}=\mu_{2}=\mu$. Consider the subspace
$\mathcal{Z}=\{z\cdot\delta_{\theta}:z\in\mathcal{Y}\}\ ,$
and let $\mathcal{P}\colon\mathcal{M}(\Theta,\mathcal{Y})\to\mathcal{Z}$ be the projection onto $\mathcal{Z}$ defined by
$\mathcal{P}\nu=\nu(\{\theta\})\cdot\delta_{\theta}\ .$
By definition of total variation, for every $\nu\in\mathcal{M}(\Theta,\mathcal{Y})$ we have
$\|\nu\|_{\operatorname{TV}}\geq\|\mathcal{P}\nu\|_{\operatorname{TV}}+\|\nu-\mathcal{P}\nu\|_{\operatorname{TV}}\ ,$
while the converse bound is simply true by triangle inequality, hence
$\|\nu\|_{\operatorname{TV}}=\|\mathcal{P}\nu\|_{\operatorname{TV}}+\|\nu-\mathcal{P}\nu\|_{\operatorname{TV}}\ .$ (15)
Note that $\mu\in\mathcal{Z}$ and thus $\mathcal{P}\mu=\mu$. Hence, applying $\mathcal{P}$ to (14) we obtain
$\mu=t\mathcal{P}\mu_{1}+(1-t)\mathcal{P}\mu_{2}\ .$ (16)
Now, consider the unit ball in $\mathcal{Z}$,
$B_{\mathcal{Z}}(1)=\{z\cdot\delta_{\theta}:z\in B_{\mathcal{Y}}(1)\}\ .$
Then $\mu\in\operatorname{Ext}(B_{\mathcal{Z}}(1))$, and $\mathcal{P}\mu_{i}\in B_{\mathcal{Z}}$ for $i=1,2$. Therefore, looking back to (16) we must have $\mathcal{P}\mu_{i}=\mu$, and in particular $\|\mathcal{P}\mu_{i}\|_{\operatorname{TV}}=\|\mu\|_{\operatorname{TV}}=1$. Moreover, $\|\mu_{i}\|_{\operatorname{TV}}\leq 1$ since $\mu_{i}\in B_{\mathcal{M}}(1)$. Thus, applying (15) to $\nu=\mu_{i}$ we get $\|\mu_{i}-\mathcal{P}\mu_{i}\|_{\operatorname{TV}}=0$, hence $\mu_{i}=\mathcal{P}\mu_{i}$, which in turn implies $\mu_{i}=\mu$, concluding the proof. ∎
The following lemma is contained in the proof of [18, Theorem 3.2]. It allows to reduce a minimization problem over compositional functions to a sequence of interpolation problems with respect to a generic minimizer. Since we found it of independent interest, we thought to emphasize it in a separate lemma.
Let $L$ be a positive integer. Take a set $\mathcal{X}_{0}$ and Banach spaces $\mathcal{X}_{1},\dots,\mathcal{X}_{L+1}$. For each $\ell=0,\ldots,L$, fix a Banach space $\mathcal{H}_{\ell}$ of functions from $\mathcal{X}_{\ell}$ to $\mathcal{X}_{\ell+1}$. To every $f{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=f_{0}\oplus\cdots\oplus f_{L}}\in\bigoplus_{\ell=0}^{L}\mathcal{H}_{\ell}$, recall that $f^{\text{deep}}:\mathcal{X}_{0}\to\mathcal{X}_{L+1}$ is defined as
$f^{\text{deep}}=f_{L}\circ\cdots\circ f_{0}\ .$
Lemma A.4.
With the above setting, fix a family $x_{1},\ldots,x_{N}\in\mathcal{X}_{0}$ and set
[…]
For $F:\mathbb{R}^{N}\to(-\infty,+\infty]$, consider the minimization problem
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\inf_{\begin{subarray}{c}f\in\bigoplus_{\ell=0}^{L}\mathcal{H}_{\ell}\end{subarray}}}F(\mathcal{A}(f))+\sum_{\ell=0}^{L}\|f_{\ell}\|_{\mathcal{H}_{\ell}}\ .$ (17)
Assume that (17) has a solution $f^{*}=\oplus f_{\ell}^{*}$, and denote $x_{i}^{(0)}=x_{i}$ and $x_{i}^{(\ell+1)}=f^{*}_{\ell}(x_{i}^{(\ell)})$ for $\ell=0,\dots,L$. Then there exists a minimizer $\widetilde{f}=\oplus_{\ell=0}^{L}\widetilde{f}_{\ell}$ of (17) such that for all $\ell=0,\dots,L$ the function $\widetilde{f}_{\ell}$ is the solution of
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\inf_{f_{\ell}\in\mathcal{H}_{\ell}}}\|f_{\ell}\|_{\mathcal{H}_{\ell}}\quad\text{subject to}\quad f_{\ell}(x^{(\ell)}_{i})=x^{(\ell+1)}_{i}\qquad i=1\dots,N\ .$ (18)
and $\|\widetilde{f}_{\ell}\|_{\mathcal{H}_{\ell}}=\|f^{*}_{\ell}\|_{\mathcal{H}_{\ell}}$.
Proof.
Let $\widetilde{f}$ be a solution to (18). The solution $f^{*}$ satisfies the constraints $f_{\ell}(x^{(\ell)}_{i})=x^{(\ell+1)}_{i}$, hence $\|\widetilde{f}_{\ell}\|_{\mathcal{H}_{\ell}}\leq\|f^{*}_{\ell}\|_{\mathcal{H}_{\ell}}$ and $\widetilde{f}_{\ell}(x_{i})=f^{*}_{\ell}(x_{i})$ for all $i=1,\dots,N$. This implies $\sum_{\ell=0}^{L}\|\widetilde{f}_{\ell}\|_{\mathcal{H}_{\ell}}\leq\sum_{\ell=0}^{L}\|f^{*}_{\ell}\|_{\mathcal{H}_{\ell}}$ and $\mathcal{A}(\widetilde{f})=\mathcal{A}(f^{*})$, so that
$F(\mathcal{A}(\widetilde{f}))+\sum_{\ell=0}^{L}\|\widetilde{f}_{\ell}\|_{\mathcal{H}_{\ell}}\leq F(\mathcal{A}(f^{*}))+\sum_{\ell=0}^{L}\|f^{*}_{\ell}\|_{\mathcal{H}_{\ell}}\ .$
[…]
The next lemmas are needed to establish continuity in the setting of our representer theorem 4.1. A continuity result akin to A.7 is also needed for the proof of [18, Theorem 3.2]. We remark that, in order for the argument to go thorough, joint continuity is required, while the proof in [18, Theorem 3.2] only establishes separate continuity. The final result remains valid since joint continuity holds true nevertheless. This can be seen as a special case of our A.7, where the last steps can be simplified in view of the fact that, in finite-dimensional spaces, weak and strong continuity coincide.
Lemma A.5.
Let $\mathcal{X}$ be a set, $\mathcal{Y}$ a Hilbert space, and $\Theta$ a locally compact, second countable topological space. Let $\rho:\mathcal{X}\times\Theta\to\mathbb{R}$ such that $\rho(x,\cdot)\in\mathcal{C}_{0}(\Theta)$ for all $x\in\mathcal{X}$, and define $\phi(x):\mathcal{M}(\Theta,\mathcal{Y})\to\mathcal{Y}$ by
$\phi(x)\mu=\int_{\Theta}\rho(x,\theta){\rm d}\mu(\theta)\qquad x\in\mathcal{X}\qquad\mu\in\mathcal{M}(\Theta,\mathcal{Y})\ .$
Then, for all $x\in\mathcal{X}$, $\phi(x)$ is continuous from $\mathcal{M}(\Theta,\mathcal{Y})$ endowed with the $*{}^{*}$ topology to $\mathcal{Y}$ endowed with the weak topology.
Proof.
Note that, since $\mathcal{Y}$ is Hilbert, the Riesz representation theorem says that $\mathcal{M}(\Theta,\mathcal{Y})=\mathcal{C}_{0}(\Theta,\mathcal{Y})^{\prime}$. Now, for all $x\in\mathcal{X}$, $\mu\in\mathcal{M}(\Theta,\mathcal{Y})$ and $y\in\mathcal{Y}$, we have
$\langle\phi(x)\mu,y\rangle_{\mathcal{Y}}={}_{\mathcal{C}_{0}(\Theta,\mathcal{Y})}\langle\rho(x,\cdot)y,\mu\rangle_{\mathcal{M}(\Theta,\mathcal{Y})}\ .$
Thus $\langle\phi(x)\cdot,y\rangle_{\mathcal{Y}}$ defines an element of the predual $\mathcal{C}_{0}(\Theta,\mathcal{Y})$, and hence it is $*{}^{*}$ continuous from $\mathcal{M}(\Theta,\mathcal{Y})$ to $\mathbb{R}$. But since this is true for all $y\in\mathcal{Y}$, it is $*{}^{*}$ continuous from $\mathcal{M}(\Theta,\mathcal{Y})$ to $\mathcal{Y}$ endowed with the weak topology. ∎
The following known result is a direct consequence of Prokhorov theorem. We report the proof for completeness.
Lemma A.6 (Joint dominated convergence theorem).
Let $\Theta$ be a Polish space. For all $n\in\mathbb{N}$, let $\lambda_{n}\in\mathcal{M}(\Theta)$ and $f_{n}:\Theta\to\mathbb{R}$ satisfying the following conditions:
  1. (i)
    for each $n\in\mathbb{N}$, the function $f_{n}$ is $\lambda_{n}$ integrable;
  2. (ii)
    the sequence $(f_{n})_{n}$ converges to some $f:\Theta\to\mathbb{R}$ uniformly on all compact sets;
  3. (iii)
    the sequence $(f_{n})_{n}$ is uniformly bounded;
  4. (iv)
    the sequence $(\lambda_{n})$ converges to some $\lambda\in\mathcal{M}(\Theta)$ with respect to the narrow topology, i.e.
    $\lim_{n\to+\infty}\int_{\Theta}\varphi(\theta){\rm d}\lambda_{n}(\theta)=\int_{\Theta}\varphi(\theta){\rm d}\lambda(\theta)\qquad\forall\varphi\in\mathcal{C}_{b}(\Theta)\ ;$
  5. (v)
    the function $f$ is $\lambda$-integrable;
Then
$\int_{\Theta}f_{n}(\theta){\rm d}\lambda_{n}(\theta)\xrightarrow[n\to\infty]{}\int_{\Theta}f(\theta){\rm d}\lambda(\theta)\ .$
Proof.
For any compact $K\subset\Theta$, we have
$\left|\int_{\Theta}f_{n}(\theta){\rm d}\lambda_{n}(\theta)-\int_{\Theta}f(\theta){\rm d}\lambda(\theta)\right|$
$=$ $\left|\int_{K}(f_{n}(\theta)-f(\theta)){\rm d}\lambda_{n}(\theta)\right|+\left|\int_{\Theta\setminus K}(f_{n}(\theta)-f(\theta)){\rm d}\lambda_{n}(\theta)\right|+\left|\int_{\Theta}(f_{n}(\theta)-f(\theta)){\rm d}\lambda(\theta)\right|$
$\leq$ $\sup_{K}|f_{n}-f||\lambda_{n}(K)|+\sup_{\Theta\setminus K}|f_{n}-f||\lambda_{n}(\Theta\setminus K)|+\left|\int_{\Theta}(f_{n}(\theta)-f(\theta)){\rm d}\lambda(\theta)\right|\ .$
The last term goes to zero by dominated convergence since $f_{n}-f$ is uniformly bounded. The first term goes to zero because $f_{n}\to f$ uniformly on $K$ and, since $(\lambda_{n})$ is convergent, $|\lambda_{n}(K)|\leq|\lambda_{n}(\Theta)|\leq\sup_{n}\|\lambda_{n}\|_{\operatorname{TV}}<\infty$. For the middle term, fix an arbitrary $\varepsilon>0$. Again, $f_{n}-f$ is uniformly bounded. Moreover, since $(\lambda_{n})$ is convergent and $\Theta$ is Polish, by the Prokhorov theorem [6, Theorem 8.6.2] there is a compact set $K_{\varepsilon}\subset\Theta$ such that $|\lambda_{n}(\Theta\setminus K_{\epsilon})|<\varepsilon$ for all $n$. By taking $K=K_{\varepsilon}$ we get
$\limsup_{n\to\infty}\left|\int_{\Theta}f_{n}(\theta){\rm d}\lambda_{n}(\theta)-\int_{\Theta}f(\theta){\rm d}\lambda(\theta)\right|\leq\left(\sup_{n}\sup_{\Theta}|f_{n}(\theta)-f(\theta)|\right)\epsilon\ .$
The claim follows because $\epsilon$ is arbitrary.
Lemma A.7.
Let $\mathcal{X}_{0}$ be a set, $\mathcal{X}_{1},\mathcal{X}_{2}$ separable Hilbert spaces, and $\Theta_{0},\Theta_{1}$ locally compact, second countable topological spaces. For $\ell=0,1$, let $\rho_{\ell}:\mathcal{X}_{\ell}\times\Theta_{\ell}\to\mathbb{R}$ such that $\rho_{\ell}(x,\cdot)\in\mathcal{C}_{0}(\Theta_{\ell})$ for all $x\in\mathcal{X}_{\ell}$, and define $\phi_{\ell}(x):\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})\to\mathcal{X}_{\ell+1}$ by
$\phi_{\ell}(x)\mu=\int_{\Theta_{\ell}}\rho_{\ell}(x,\theta){\rm d}\mu(\theta)\qquad x\in\mathcal{X}_{\ell}\qquad\mu\in\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})\ .$
Assume there are $C>0$, $g\in\mathcal{C}_{b}(\Theta_{1},\mathcal{X}_{1})$ and $\beta\in\mathcal{C}_{0}(\Theta_{1})$ such that, for all $x,x^{\prime}\in\mathcal{X}_{1}$ and $\theta\in\Theta_{1}$,
$|\rho_{1}(x,\theta)-\rho_{1}(x^{\prime},\theta)|\leq C|\langle x-x^{\prime},g(\theta)\rangle_{\mathcal{X}_{1}}||\beta(\theta)|\ .$ (19)
Let $r_{0},r_{1}>0$. Then, for all $x\in\mathcal{X}_{0}$, the map
$\Gamma_{x}(\mu,\nu)=\phi_{1}((\phi_{0}(x)\mu))\nu$
is jointly $*{}^{*}$ continuous from $B_{\mathcal{M}(\Theta_{0},\mathcal{X}_{1})}(r_{0})\times B_{\mathcal{M}(\Theta_{1},\mathcal{X}_{2})}(r_{1})$ to $\mathcal{X}_{2}$ endowed with the weak topology.
Proof.
By the Banach–Alaoglu theorem, the product $B=B_{\mathcal{M}(\Theta_{0},\mathcal{X}_{1})}(r_{0})\times B_{\mathcal{M}(\Theta_{1},\mathcal{X}_{2})}(r_{1})$ is compact. Moreover, since $\mathcal{X}_{\ell+1}$ ($\ell=0,1$) is separable, so is $\mathcal{C}_{0}(\Theta_{\ell},\mathcal{X}_{\ell+1})$ by the Stone–Weierstrass theorem. Also, $\mathcal{X}_{\ell+1}$ is Hilbert, hence $\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})=\mathcal{C}_{0}(\Theta_{\ell},\mathcal{X}_{\ell+1})^{\prime}$ by the Riesz representation theorem. Therefore, $B$ is metrizable [20, Theorem 3.16]. Thus, it is enough to prove the ($*{}^{*}$-weak) sequential continuity of $\Gamma_{x}$.
To this end, let $(\mu_{n},\nu_{n})\to(\mu,\nu)$ ($*{}^{*}$). We want to show that $\Gamma_{x}(\mu_{n},\nu_{n})\to\Gamma_{x}(\mu,\nu)$ (weakly). We have
$\Gamma_{x}(\mu_{n},\nu_{n})-\Gamma_{x}(\mu,\nu)=[\phi_{1}(\phi_{0}(x)\mu_{n})-\phi_{1}(\phi_{0}(x)\mu)]\nu_{n}+\phi_{1}(\phi_{0}(x)\mu)(\nu_{n}-\nu)\ .$
The second term goes to zero by A.5. Let us call $\mathcal{I}$ the first term, and let $z_{n}=\phi_{0}(x)\mu_{n}$, $z=\phi_{0}(x)\mu$. For all $y\in\mathcal{X}_{2}$, by assumption (19) we have
$|\langle\mathcal{I},y\rangle_{\mathcal{X}_{2}}|$ $\leq\int_{\Theta_{1}}|\rho_{1}(z_{n},\theta)-\rho_{1}(z,\theta)||\beta(\theta)|{\rm d}|[\nu_{n}]_{y}|(\theta)$
$\leq C\int_{\Theta_{1}}|\langle z_{n}-z,g(\theta)\rangle_{\mathcal{X}_{1}}||\beta(\theta)|{\rm d}|[\nu_{n}]_{y}|(\theta)\ ,$
where $[\nu_{n}]_{y}(E)=\langle\nu_{n}(E),y\rangle_{\mathcal{X}_{2}}$ for all $E\in\mathcal{B}(\Theta_{1})$. We want to apply A.6 with $f_{n}(\theta)=|\langle z_{n}-z,g(\theta)\rangle_{\mathcal{X}_{1}}|$ and $\lambda_{n}=|\beta||[\nu_{n}]_{y}|$, so to conclude that $|\langle\mathcal{I},y\rangle_{\mathcal{X}_{2}}|\to 0$.
First, let us verify that $(f_{n})$ is uniformly bounded. We have
$f_{n}(\theta)\leq\|z_{n}-z\|_{\mathcal{X}_{1}}\|g(\theta)\|_{\mathcal{X}_{1}}\leq\sup_{n}\|z_{n}-z\|_{\mathcal{X}_{1}}\|g\|_{\infty}\ ,$
where $\sup_{n}\|z_{n}-z\|_{\mathcal{X}_{1}}<\infty$ because $(z_{n}-z)$ is convergent, and $\|g\|_{\infty}<\infty$ by assumption. Next, we show that $(\lambda_{n})$ converges pointwise on $\mathcal{C}_{b}(\Theta_{1})$. Let $\lambda=|\beta||[\nu]_{y}|$ where $[\nu]_{y}(E)=\langle\nu(E),y\rangle_{\mathcal{X}_{2}}$ for all $E\in\mathcal{B}(\Theta_{1})$. Then, for every $h\in\mathcal{C}_{b}(\Theta_{1})$ we have $h|\beta|\in\mathcal{C}_{0}(\Theta_{1})$, and hence, since $[\nu_{n}]_{y}\to[\nu]_{y}$ pointwise on $\mathcal{C}_{0}(\Theta_{1})$,
$\int_{\Theta_{1}}h(\theta){\rm d}\lambda_{n}(\theta)=\int_{\Theta_{1}}h(\theta)|\beta(\theta)|{\rm d}[\nu_{n}]_{y}(\theta)\to\int_{\Theta_{1}}h(\theta)|\beta(\theta)|{\rm d}[\nu]_{y}(\theta)=\int_{\Theta_{1}}h(\theta){\rm d}\lambda(\theta)\ .$
Finally, we show that $f_{n}\to 0$ uniformly on compact sets. Let $K\subset\Theta_{1}$ be compact, and fix an arbitrary $\varepsilon>0$. Since $g$ is continuous, $g(K)$ is compact in $\mathcal{X}_{1}$, and thus it can be covered by a finite number of closed balls. Let $w_{1},\dots,w_{q}\in\mathcal{X}_{1}$ be the centers of such balls, and define $P:\mathcal{X}_{1}\to\mathcal{X}_{1}$ as the projection onto $\operatorname{span}\{w_{1},\dots,w_{q}\}$. Then $\sup_{w\in g(K)}\|w-Pw\|_{\mathcal{X}_{!}}\leq\varepsilon$. Hence, for every $\theta\in K$ there is $w\in P\mathcal{X}_{1}$ such that $\|g(\theta)-w\|_{\mathcal{X}_{1}}\leq\varepsilon$. Thus, we have
$|f_{n}(\theta)|$ $=|\langle z_{n}-z,g(\theta)\rangle_{\mathcal{X}_{1}}|$
$\leq|\langle z_{n}-z,g(\theta)-w\rangle_{\mathcal{X}_{1}}|+|\langle P(z_{n}-z),w\rangle_{\mathcal{X}_{1}}|$
$\leq\|z_{n}-z\|_{\mathcal{X}_{1}}\|g(\theta)-w\|_{\mathcal{X}_{1}}+\|P(z_{n}-z)\|_{\mathcal{X}_{1}}\|w\|_{\mathcal{X}_{1}}$
$\leq\|z_{n}-z\|_{\mathcal{X}_{1}}\|g(\theta)-w\|_{\mathcal{X}_{1}}+\|P(z_{n}-z)\|_{\mathcal{X}_{1}}\left(\|g(\theta)-w\|_{\mathcal{X}_{1}}+\|g(\theta)\|_{\mathcal{X}_{1}}\right)$
$\leq\|z_{n}-z\|_{\mathcal{X}_{1}}\varepsilon+\|P(z_{n}-z)\|_{\mathcal{X}_{1}}\left(\varepsilon+\|g\|_{\infty}\right)\ .$
Now, since $z_{n}\to z$ weakly, $\sup_{n}\|z_{n}-z\|_{\mathcal{X}_{1}}<\infty$, and $\|P(z_{n}-z)\|_{\mathcal{X}_{1}}\to 0$ because $P$ has finite rank.
All the assumptions of A.6 are therefore satisfied, and its application concludes the proof. ∎